# Towards Integrating Emerging AI Applications in SE Education

### v1

@@ 2.2 Related Work
Several systematic reviews have investigated and classified aspects of AI applications in education. Bittencourt et al. [bittencourt2023positive] conducted an analysis exploring the intersection of positive psychology and AI in education (P-AIED), identifying P-AIED as a new global movement focusing on positive emotion and engagement related to AI-supported teaching and learning. Kuhail et al. [kuhail2023interacting] and Okonkwo and Ade-Ibijola [25] performed studies related to the application of chatbots in education. Kuhail et al. analyzed work on educational chatbots and found that these are primarily used in computer science, language learning, general education, and to a lesser extent in fields like engineering and mathematics. Most chatbots were web-based and more than half functioned as teaching agents, while over a third acted as peer agents. Okonkwo and Ade-Ibijola analyze 53 primary studies and identify an increasing trend in the application of chatbots in education.
Zhai et al. [zhai2021review] analyzed 100 papers focusing on the application of AI in the education sector from 2010 to 2020. They classified primary studies into a development layer (classification, matching, recommendation, and deep learning), an application layer (feedback, reasoning, and adaptive learning), and an integration layer (affection computing, role-playing, immersive learning, and gamification). Chen et al. [chen2020] performed a study investigating the impact of AI on education, focusing on its application and effects in administrative, instructional, and learning contexts. Similarly, Zawacki-Richter et al. [zawacki2019systematic] conducted a systematic review of research on the application of AI in higher education. The majority of research comes from the fields of Computer Science and STEM with AI being mainly used in areas of student profiling and predictive analytics, as well as intelligent tutoring systems.
Several studies have focused on AI in computer science education. Denny et al. [denny2023computing] explore the challenges and opportunities presented by recent advances in AI, particularly code generation models, and their potential impact on computing education. Kalles [16] reports on the experience of using AI systems as the basis of educating IT students. The application areas were mainly decision tree lifecycle management and board game learning mechanisms. Daun and Brings [6] discuss the role of generative AI technologies, such as ChatGPT, in SE education. They discuss potential risks associated with students’ use of generative AI but also identify several opportunities that these technologies can offer in enhancing educational practices such as individualization of education and personalized feedback.

### v2

@@ 2.2 Related Work
Several systematic reviews have investigated and classified aspects of AI applications in education. Bittencourt et al. [4] conducted an analysis exploring the intersection of positive psychology and AI in education (P-AIED), identifying P-AIED as a new global movement focusing on positive emotion and engagement related to AI-supported teaching and learning. Kuhail et al. [21] and Okonkwo and Ade-Ibijola [29] performed studies related to the application of chatbots in education. Kuhail et al. analyzed work on educational chatbots and found that these are primarily used in computer science, language learning, general education, and to a lesser extent in fields like engineering and mathematics. Most chatbots were web-based and more than half functioned as teaching agents, while over a third acted as peer agents. Okonkwo and Ade-Ibijola analyze 53 primary studies and identify an increasing trend in the application of chatbots in education.
Zhai et al. [41] analyzed 100 papers focusing on the application of AI in the education sector from 2010 to 2020. They classified primary studies into a development layer (classification, matching, recommendation, and deep learning), an application layer (feedback, reasoning, and adaptive learning), and an integration layer (affection computing, role-playing, immersive learning, and gamification). Chen et al. [6] performed a study investigating the impact of AI on education, focusing on its application and effects in administrative, instructional, and learning contexts. Similarly, Zawacki-Richter et al. [40] conducted a systematic review of research on the application of AI in higher education. The majority of research comes from the fields of Computer Science and STEM with AI being mainly used in areas of student profiling and predictive analytics, as well as intelligent tutoring systems.
Several studies have focused on AI in computer science education. Denny et al. [9] explore the challenges and opportunities presented by recent advances in AI, particularly code generation models, and their potential impact on computing education. Kalles [19] reports on the experience of using AI systems as the basis of educating IT students. The application areas were mainly decision tree lifecycle management and board game learning mechanisms. Daun and Brings [8] discuss the role of generative AI technologies, such as ChatGPT, in SE education. They discuss potential risks associated with students’ use of generative AI but also identify several opportunities that these technologies can offer in enhancing educational practices such as individualization of education and personalized feedback.